# Off-Policy Evaluation of Slate Bandit Policies via Optimizing Abstraction

### 03 Feb 2024 v1

@@ Abstract.
journalyear: 2024conference: Woodstock ’24: ACM Symposium on Neural Gaze Detection; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 3. The LIPS Estimator
where $\pi(z|\bm{x}):=\sum_{\bm{s}\in\{\bm{s}^{\prime}\in\mathcal{S}\mid\phi_{\theta}(\bm{s}^{\prime})=z\}}\pi(\bm{s}|\bm{x})$ is a marginal distribution of an abstracted slate induced by policy $\pi$, and $w(\bm{x},z):=\pi(z|\bm{x})/\pi_{0}(z|\bm{x})$ is the latent importance weight. Note that we can readily extend LIPS to the case with a context-dependent and stochastic abstraction based on a parameterized distribution $p_{\theta}:\mathcal{X}\times\mathcal{S}\rightarrow\Delta(\mathcal{Z})$. This means that we can generalize Eq. (3) as
@@
where $z_{i}\sim p_{\theta}(\cdot\,|\,\bm{x}_{i},\bm{s}_{i})$ and $p_{\theta}(z\,|\,\bm{x};\pi):=\sum_{\bm{s}\in\mathcal{S}}\pi(\bm{s}\,|\,\bm{x})p_{\theta}(z\,|\,\bm{x},\bm{s})$ $=\sum_{\bm{s}\in\mathcal{S}}p_{\theta}(\bm{s},z\,|\,\bm{x};\pi)$. This extension allows for a flexible control of the bias-variance tradeoff of LIPS and ensures a tractable optimization of abstraction. The central idea of LIPS is to circumvent the reliance on slate- or slot-wise importance weights, substantially improving the variance from IPS and PI while avoiding the linear-reward assumption like PI. The following formally analyzes LIPS and shows its statistical advantages over the existing estimators. We also develop a data-driven optimization procedure for slate abstractions to directly minimize the MSE of LIPS.
@@ 3.1. Theoretical Analysis
An abstraction function is sufficient if it aggregates only slates that have the same expected reward, and it means that the latent slate space $\mathcal{Z}$ retains sufficient information to characterize the reward function. For example, identity abstraction $\phi(\bm{s})=\bm{s}$ is always sufficient (LIPS is reduced to IPS in this case). Note here that this notion of sufficiency does not impose any particular restriction on the reward function form such as linearity. Furthermore, sufficient slate abstractions may not be unique, and there could potentially be many sufficient slate abstractions.
@@
When the slate $\bm{s}$ is almost deterministic given $(\bm{x},z)$, the value of $p_{\theta}(\bm{s}_{j}|\bm{x},z;\pi_{0})$ tends to be either zero or one. As a result, the product $p_{\theta}(\bm{s}_{j}|\bm{x},z;\pi_{0})p_{\theta}(\bm{s}_{k}|\bm{x},z;\pi_{0})$ (the first factor) becomes nearly zero. This indicates that if the latent slate space sufficiently encapsulates the information required to reconstruct the original slates, then even with a stochastic slate abstraction, the bias of LIPS remains minimal. Furthermore, the second factor relates to the predictiveness of rewards given the latent variable $z$. Specifically, if the latent variable is almost sufficient—meaning that slates with similar expected rewards tend to have similar abstraction distributions—the difference $q(\bm{x},\bm{s}_{j})-q(\bm{x},\bm{s}_{k})$ will be small when $p_{\theta}(\bm{s}_{j}|\bm{x},z;\pi_{0})p_{\theta}(\bm{s}_{k}|\bm{x},z;\pi_{0})$ is large, leading to a reduced bias of LIPS. Thus, the analysis implies that the bias of LIPS remains small if the latent slate space is finer-grained and abstraction is closer to deterministic, which makes the latent variable $z$ more predictive of slate $\bm{s}$ and reward $r$.
@@
There are two primary factors that determine the degree of variance reduction. The first is the second moment of the reward, which increases when the reward has high variability. The second is the variance of the slate-wise importance weight $w(\bm{x},\bm{s})$ relative to the conditional distribution $\pi_{0}(\bm{s}|\bm{x},\phi(\bm{s})),(=\pi_{0}(\bm{s}|\bm{x})/\pi_{0}(\phi(\bm{s})|\bm{x}))$. This factor escalates when (i) $w(\bm{x},\bm{s})$ attains large values and (ii) $\pi_{0}(\bm{s}|x,\phi(\bm{s}))$ maintains sufficient stochasticity. This suggests that variance reduction of LIPS could be exponential, as the variance of the slate importance weight may increase exponentially with the slate size $L$. Additionally, the degree of variance reduction can be modulated through the cardinality of the latent slate space and the entropy of the slate abstraction. In other words, a more generalized and stochastic abstraction leads to greater variance reduction. A key theoretical insight here is that a sufficient abstraction might not always minimize the MSE of LIPS. Table 1 presents a hypothetical example demonstrating how LIPS with an insufficient abstraction may achieve a lower MSE than with a sufficient abstraction. Specifically, LIPS with a sufficient abstraction achieves the MSE of 1.0 with zero bias as per Theorem 3.2. However, a lower MSE can be realized by intentionally using an insufficient abstraction. This approach allows for a substantial reduction in variance (-0.8) at the cost of a minor increase in squared bias (+0.25). Thus, a sufficient abstraction does not necessarily yield the optimal MSE for LIPS. Consequently, rather than seeking a sufficient abstraction, the following sections describe a data-driven strategy to optimize the abstraction to minimize the MSE of LIPS.
@@ 3.2. Optimizing Slate Abstractions
$\theta$, $\psi$, and $\omega$ are the parameters of slate abstraction $p_{\theta}(z|\bm{x},\bm{s};\pi_{0})$, slate reconstruction $p_{\psi}(\bm{s}|\bm{x},z;\pi_{0})$, and reward construction $\hat{q}_{\omega}(\bm{x},z)$ models. [^2] The pseudo-code (Algorithm 1) of this abstraction optimization procedure can be found in the appendix. Appendix B provides a more extensive bias-variance analysis of LIPS when given a stochastic abstraction $p_{\theta}$. The first two terms of the loss function aim for bias reduction, while the last term controls variance reduction of LIPS. More specifically, the first term measures the identifiability of the slates and the second term measures how predictive the rewards are based on the latent variable. The final term acts as a regulator for the stochasticity of $p_{\theta}(\bm{s}|\bm{x},z;\pi_{0})$ by aligning $p_{\theta}(z,|,\bm{x},\bm{s};\pi_{0})$ more closely with $p_{\psi}(z,|,\bm{x};\pi_{0})$. This will indeed regularize the latent importance weights and make them close to $1$ everywhere (as described in the appendix in detail), and thus we can expect a larger variance reduction of LIPS when the last term is dominant. The regularization weight $\beta$ is thus considered the key hyper-parameter that governs the bias-variance tradeoff of LIPS (i.e., a smaller $\beta$ implies a smaller bias and larger variance of LIPS, while a larger $\beta$ leads to a larger bias and smaller variance). This hyper-parameter can be tuned by existing parameter tuning methods such as SLOPE (Su et al., 2020b) and PAS-IF (Udagawa et al., 2023), which are feasible using only observed logged data $\mathcal{D}$. Section 5 empirically investigates how LIPS works with these existing parameter tuning methods.
@@ 4. Related Work
Figure 1. Comparison of the estimators’ MSE (normalized by the true value $V(\pi)$) with varying slate sizes ($L$) and with reward functions (1) - (3) on the Wiki10-31K dataset.
Figure 2. Comparison of the estimators’ MSE (normalized by the true value $V(\pi)$) with varying slate sizes ($L$) and with reward functions (1) - (3) on the Eurlex-4K dataset.
Figure 3. Comparison of the estimators’ MSE (normalized by the true value $V(\pi)$) with varying data sizes ($n$) and with reward functions (1) - (3) on the Wiki10-31K dataset.
Figure 4. Comparison of the estimators’ MSE (normalized by the true value $V(\pi)$) with varying data sizes ($n$) and with reward functions (1) - (3) on the Eurlex-4K dataset.
@@ 5. Empirical Evaluation
This section empirically compares LIPS with many relevant estimators on two real-world datasets, namely Wiki10-31K and Eurlex-4K provided in the Extreme Classification repository (Bhatia et al., 2016). Note that our experiment code is available at https://github.com/aiueola/webconf2024-slate-ope-via-abstraction.
@@ 5.1. Experiment Setup
We follow the standard ”supervised-to-bandit” procedure to conduct an OPE experiment based on classification data, as used in many previous studies (Farajtabar et al., 2018; Su et al., 2020a, 2019; Wang et al., 2017). Specifically, we use the extreme classification datasets called Wiki10-31K and Eurlex-4K. These datasets consist of many documents associated with a large number of labels, which are approximately 31K for Wiki10-31K and 4K for Eurlex-4K. More detailed dataset statistics are provided in Appendix A.
@@
We compare LIPS with Direct Method (DM) (Beygelzimer and Langford, 2009), IPS (Strehl et al., 2010), PI (Swaminathan et al., 2017), and MIPS (Saito and Joachims, 2022b). DM is a regression-based estimator that estimates the policy value based on an estimated reward function $\hat{q}(\bm{x},\bm{s})$, which is learned by a neural network in our experiment. For LIPS, we employ a discrete abstraction whose dimension is 100 ($|\mathcal{Z}|=100$). When optimizing a slate abstraction, we select the hyper-parameter $\beta$ from $\{0.01,0.1,1.0,10.0\}$ based only on the available logged data via the SLOPE algorithm (Su et al., 2020b; Tucker and Lee, 2021). Appendix A provides some more details of SLOPE and describes how we parameterize and optimize the slate abstraction distribution $p_{\theta}$. Note that MIPS defines its importance weight taking only relevant slots (i.e., $l=1,2,\ldots,\lfloor L/2\rfloor$) into account and is defined as $\hat{V}_{\mathrm{MIPS}}(\pi;\mathcal{D}):=\frac{1}{n}\sum_{i=1}^{n}\big{(}\prod_{l=1}^{\lfloor L/2\rfloor}\frac{\pi(a_{i,l}\,|\,\bm{x}_{i})}{\pi_{0}(a_{i,l}\,|\,\bm{x}_{i})}\big{)}r_{i},$ where the importance weights use only the first $\lfloor L/2\rfloor$ slots as an action embedding leveraging the fact that the reward functions (1) - (3) depend only on these slots. MIPS is unbiased and has a lower variance than IPS, however, it is infeasible in practice since we do not know the true reward function. We include MIPS in our experiments since it is useful to investigate the effectiveness of intentionally using an insufficient abstraction.
In addition to these baselines, we also report the results of “LIPS (w/ best $\beta$)” as a reference. It indicates LIPS with the best value of $\beta$ selected based on the ground-truth MSE, which provides the best accuracy achievable by our LIPS framework with an oracle hyper-parameter selection.
Figure 5. Comparison between LIPS and the DR estimators wrt their MSE (normalized by the true value $V(\pi)$) under varying slate sizes ($L$) and with reward functions (1) - (3) on the Wiki10-31K dataset.
Figure 6. Comparison between LIPS and the DR estimators wrt their MSE (normalized by the true value $V(\pi)$) under varying slate sizes ($L$) and with reward functions (1) - (3) on the Eurlex-4K dataset.
Figure 7. Comparison between LIPS and the DR estimators wrt their MSE (normalized by the true value $V(\pi)$) under varying data sizes ($n$) and with reward functions (1) - (3) on the Wiki10-31K dataset.
Figure 8. Comparison between LIPS and the DR estimators wrt their MSE (normalized by the true value $V(\pi)$) under varying data sizes ($n$) and with reward functions (1) - (3) on the Eurlex-4K dataset.
@@ 5.2. Results and Discussion
The following discusses the results obtained by running OPE simulations with 50 different logged datasets generated under different random seeds. We compare the estimators’ accuracy by their MSE normalized by the true policy value of the target policy, which is defined as $\mathrm{MSE}(\hat{V}(\pi))/V(\pi)$. We use $L=8$, $|\mathcal{A}_{l}|=10\,(\forall l\in[L])$, and $n=4000$ as default experiment parameters.
@@
Beyond the basic comparisons against DM, IPS, PI, and MIPS, we finally compare LIPS with a set of hybrid estimators including DR (Dudík et al., 2011), PI-DR (Su et al., 2020a), and OffCEM (Saito et al., 2023) (Appendix A defines these estimators in detail with math notations). Figures 8 and 8 show this comparison on Wiki10-31K and Eurlex-4K for varying slate sizes $(L\in\{4,6,\ldots,12\})$. Figures 8 and 8 compare the methods on the datasets for varying data sizes ($n\in\{1000,2000,\ldots,16000\}$). The results demonstrate that LIPS outperforms these hybrid estimators for a range of experiment values ($L$ and $n$) and non-linear reward functions. This is because DR has the variance issue due to its slate-wise importance weighting while PI-DR suffers from bias due to its linearity assumption and variance due to its slot-wise importance weighting. OffCEM is an extension of MIPS and uses a sufficient abstraction, but it is not optimized towards the MSE, and thus produces much larger variance than LIPS. Note that we also observe the similar trend on an additional dataset as reported in Appendix A. These empirical observations suggest that, for OPE of slate bandits, tuning the definition of importance weights by optimizing slate abstractions with our method is more crucial and effective than adding a reward estimator ($\hat{q}$) as in the hybrid estimators.

### 17 Feb 2024 v2

@@ Abstract.
off-policy evaluation, slate bandits, inverse propensity score.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singaporebooktitle: Proceedings of the ACM Web Conference 2024 (WWW ’24), May 13–17, 2024, Singapore, Singaporedoi: 10.1145/3589334.3645343isbn: 979-8-4007-0171-9/24/05ccs: Information systems Recommender systemsccs: Information systems Evaluation of retrieval results
@@ 3. The LIPS Estimator
where $\pi(z|\bm{x}):=\sum_{\bm{s}\in\mathcal{S}}\mathbb{I}\{\phi_{\theta}(\bm{s})=z\}\pi(\bm{s}|\bm{x})$ is a marginal distribution of an abstracted slate induced by policy $\pi$, and $w(\bm{x},z):=\pi(z|\bm{x})/\pi_{0}(z|\bm{x})$ is the latent importance weight. Note that we can readily extend LIPS to the case with a context-dependent and stochastic abstraction based on a parameterized distribution $p_{\theta}:\mathcal{X}\times\mathcal{S}\rightarrow\Delta(\mathcal{Z})$. This means that we can generalize Eq. (3) as
@@
where $z_{i}\sim p_{\theta}(z\,|\,\bm{x}_{i},\bm{s}_{i})$ and $p_{\theta}(z\,|\,\bm{x};\pi):=\sum_{\bm{s}\in\mathcal{S}}\pi(\bm{s}\,|\,\bm{x})p_{\theta}(z\,|\,\bm{x},\bm{s})$ $=\sum_{\bm{s}\in\mathcal{S}}p_{\theta}(\bm{s},z\,|\,\bm{x};\pi)$. This extension allows for a flexible control of the bias-variance tradeoff of LIPS and ensures a tractable optimization of abstraction. The central idea of LIPS is to circumvent the reliance on slate- or slot-wise importance weights, substantially improving the variance from IPS and PI while avoiding the linear-reward assumption like PI. The following formally analyzes LIPS and shows its statistical advantages over the existing estimators. We also develop a data-driven optimization procedure for slate abstractions to directly minimize the MSE of LIPS.
@@ 3.1. Theoretical Analysis
An abstraction function is sufficient if it aggregates only the slates that have the same expected reward, and it means that the latent slate space $\mathcal{Z}$ retains sufficient information to characterize the reward function. For example, identity abstraction $\phi(\bm{s})=\bm{s}$ is always sufficient (LIPS is reduced to IPS in this case). Note here that this notion of sufficiency does not impose any particular restriction on the reward function form such as linearity. Furthermore, sufficient slate abstractions may not be unique, and there could potentially be many sufficient slate abstractions.
@@
When the slate $\bm{s}$ is almost deterministic given $(\bm{x},z)$, the value of $p_{\theta}(\bm{s}_{j}|\bm{x},z;\pi_{0})$ tends to be either zero or one. As a result, the product $p_{\theta}(\bm{s}_{j}|\bm{x},z;\pi_{0})p_{\theta}(\bm{s}_{k}|\bm{x},z;\pi_{0})$ becomes nearly zero. This indicates that if the latent slate space sufficiently encapsulates the information required to reconstruct the original slates, then even with a stochastic slate abstraction, the bias of LIPS remains minimal. Furthermore, the second factor relates to the predictiveness of rewards given the latent variable $z$. Specifically, if the latent variable is almost sufficient—meaning that slates with similar expected rewards tend to have similar abstraction distributions—the difference $q(\bm{x},\bm{s}_{j})-q(\bm{x},\bm{s}_{k})$ will be small, leading to a reduced bias of LIPS. Thus, the analysis implies that the bias of LIPS remains small if the latent slate space is finer-grained and abstraction is closer to deterministic, which makes the latent variable $z$ more predictive of slate $\bm{s}$ and reward $r$.
@@
There are two primary factors that determine the degree of variance reduction. The first is the second moment of the reward, which increases when the reward has high variability. The second is the variance of the slate-wise importance weight $w(\bm{x},\bm{s})$ regarding the conditional distribution $\pi_{0}(\bm{s}|\bm{x},\phi(\bm{s}))(=\pi_{0}(\bm{s}|\bm{x})/\pi_{0}(\phi(\bm{s})|\bm{x}))$. This factor escalates when (i) $w(\bm{x},\bm{s})$ attains large values and (ii) $\pi_{0}(\bm{s}|x,\phi(\bm{s}))$ maintains sufficient stochasticity. This suggests that variance reduction of LIPS could be exponential, as the variance of the slate importance weight may increase exponentially with the slate size $L$. Additionally, the degree of variance reduction can be modulated through the cardinality of the latent slate space and the entropy of the slate abstraction. In other words, a coarser and stochastic abstraction leads to greater variance reduction. A key theoretical insight here is that a sufficient abstraction might not always minimize the MSE of LIPS. Table 1 presents a hypothetical example demonstrating how LIPS with an insufficient abstraction may achieve a lower MSE than with a sufficient abstraction. Specifically, LIPS with a sufficient abstraction achieves the MSE of 1.0 with zero bias as per Theorem 3.2. However, a lower MSE can be realized by intentionally using an insufficient abstraction. This approach allows for a substantial reduction in variance (-0.8) at the cost of a minor increase in squared bias (+0.25). Thus, a sufficient abstraction does not necessarily yield the optimal MSE for LIPS. Consequently, rather than seeking a sufficient abstraction, the following sections describe a data-driven strategy to optimize the abstraction to minimize the MSE of LIPS.
@@ 3.2. Optimizing Slate Abstractions
$\theta$, $\psi$, and $\omega$ are the parameters of slate abstraction $p_{\theta}(z\,|\,\bm{x},\bm{s};\pi_{0})$, slate reconstruction $p_{\psi}(\bm{s}\,|\,\bm{x},z;\pi_{0})$, and reward construction $\hat{q}_{\omega}(\bm{x},z)$ models. [^2] The pseudo-code (Algorithm 1) of this abstraction optimization procedure can be found in the appendix. Appendix B provides a more extensive bias-variance analysis of LIPS when given a stochastic abstraction $p_{\theta}$. Note that, if some action features are available, it is also possible to use such action features concatenated across slates to the models to further improve the abstraction optimization procedure. The first two terms of the loss function aim for bias reduction, while the last term controls variance reduction of LIPS. More specifically, the first term measures the identifiability of the slates and the second term measures how predictive the rewards are based on the latent variable. The final term acts as a regulator for the stochasticity of $p_{\theta}(\bm{s}\,|\,\bm{x},z;\pi_{0})$ by aligning $p_{\theta}(z\,|\,\bm{x},\bm{s};\pi_{0})$ more closely with $p_{\psi}(z\,|\,\bm{x};\pi_{0})$. This will indeed regularize the latent importance weights and make them close to $1$ everywhere (as described in the appendix in detail), and thus we can expect a larger variance reduction of LIPS when the last term is dominant. The regularization weight $\beta$ is thus considered the key hyper-parameter that governs the bias-variance tradeoff of LIPS (i.e., a smaller $\beta$ implies a smaller bias and larger variance of LIPS, while a larger $\beta$ leads to a larger bias and smaller variance). This hyper-parameter can be tuned by existing parameter tuning methods such as SLOPE (Su et al., 2020b) and PAS-IF (Udagawa et al., 2023), which are feasible using only observed logged data $\mathcal{D}$. Section 5 empirically investigates how LIPS works with these existing parameter tuning methods.
@@ 4. Related Work
Figure 1. Comparison of the estimators’ MSE (normalized by the true value $V(\pi)$) with varying slate sizes ($L$) and with reward functions (1) - (3) on the Wiki10-31K dataset. Note that the y-axis is on a log scale, thus the fluctuations observed for LIPS’s MSE are less than 0.01.
Figure 2. Comparison of the estimators’ MSE (normalized by the true value $V(\pi)$) with varying slate sizes ($L$) and with reward functions (1) - (3) on the Eurlex-4K dataset. Note that the y-axis is on a log scale, thus the fluctuations observed for LIPS’s MSE are less than 0.01.
Figure 3. Comparison of the estimators’ MSE (normalized by the true value $V(\pi)$) with varying data sizes ($n$) and with reward functions (1) - (3) on the Wiki10-31K dataset. Note that the y-axis is on a log scale.
Figure 4. Comparison of the estimators’ MSE (normalized by the true value $V(\pi)$) with varying data sizes ($n$) and with reward functions (1) - (3) on the Eurlex-4K dataset. Note that the y-axis is on a log scale.
@@ 5. Empirical Evaluation
This section empirically compares LIPS with many relevant estimators on two real-world datasets. Our experiment code is available at https://github.com/aiueola/webconf2024-slate-ope-via-abstraction.
@@ 5.1. Experiment Setup
We follow the standard ”supervised-to-bandit” procedure to conduct an OPE experiment based on classification data, as used in many previous studies (Farajtabar et al., 2018; Su et al., 2020a, 2019; Wang et al., 2017). Specifically, we use the extreme classification datasets called Wiki10-31K and Eurlex-4K (Bhatia et al., 2016). These datasets consist of many documents associated with a large number of labels, which are approximately 31K for Wiki10-31K and 4K for Eurlex-4K. Detailed dataset statistics are provided in Appendix A.
@@
We compare LIPS with the Direct Method (DM) (Beygelzimer and Langford, 2009), IPS (Strehl et al., 2010), PI (Swaminathan et al., 2017), and MIPS (Saito and Joachims, 2022b). DM is a regression-based estimator that estimates the policy value based on an estimated reward function $\hat{q}(\bm{x},\bm{s})$, which is learned by a neural network in our experiment. For LIPS, we employ a discrete abstraction whose dimension is 100 ($|\mathcal{Z}|=100$). When optimizing a slate abstraction, we select the hyper-parameter $\beta$ from $\{0.01,0.1,1.0,10.0\}$ based only on the available logged data via the SLOPE algorithm (Su et al., 2020b; Tucker and Lee, 2021). Appendix A provides some more details of SLOPE and describes how we parameterize and optimize the slate abstraction distribution $p_{\theta}$. Note that MIPS defines its importance weight taking only relevant slots (i.e., $l=1,2,\ldots,\lfloor L/2\rfloor$) into account and is defined as $\hat{V}_{\mathrm{MIPS}}(\pi;\mathcal{D}):=\frac{1}{n}\sum_{i=1}^{n}\big{(}\prod_{l=1}^{\lfloor L/2\rfloor}\frac{\pi(a_{i,l}\,|\,\bm{x}_{i})}{\pi_{0}(a_{i,l}\,|\,\bm{x}_{i})}\big{)}r_{i},$ where the importance weights use only the first $\lfloor L/2\rfloor$ slots as an action embedding leveraging the fact that the reward functions (1) - (3) depend only on these slots. MIPS is unbiased and has a lower variance than IPS, however, it is infeasible in practice since we do not know the true reward function. We include MIPS in our experiments since it is useful to investigate the effectiveness of intentionally using an insufficient abstraction. In addition to these baselines, we also report the results of “LIPS (w/ best $\beta$)” as a reference. It indicates LIPS with the best value of $\beta$ selected based on the ground-truth MSE, which provides the best accuracy achievable by our LIPS framework with an oracle hyper-parameter selection.
Figure 5. Comparison between LIPS and the DR estimators wrt their MSE (normalized by the true value $V(\pi)$) under varying slate sizes ($L$) and with reward functions (1) - (3) on the Wiki10-31K dataset. Note that the y-axis is on a log scale.
Figure 6. Comparison between LIPS and the DR estimators wrt their MSE (normalized by the true value $V(\pi)$) under varying slate sizes ($L$) and with reward functions (1) - (3) on the Eurlex-4K dataset. Note that the y-axis is on a log scale.
Figure 7. Comparison between LIPS and the DR estimators wrt their MSE (normalized by the true value $V(\pi)$) under varying data sizes ($n$) and with reward functions (1) - (3) on the Wiki10-31K dataset. Note that the y-axis is on a log scale.
Figure 8. Comparison between LIPS and the DR estimators wrt their MSE (normalized by the true value $V(\pi)$) under varying data sizes ($n$) and with reward functions (1) - (3) on the Eurlex-4K dataset. Note that the y-axis is on a log scale.
@@ 5.2. Results and Discussion
The following discusses the results obtained by running OPE simulations with 50 different logged datasets generated under different random seeds. We compare the estimators’ accuracy by their MSE normalized by the true policy value of the target policy, which is defined as $\mathrm{MSE}(\hat{V}(\pi))/(V(\pi))^{2}$. We use $L=8$, $|\mathcal{A}_{l}|=10\,(\forall l\in[L])$, and $n=4000$ as default experiment parameters.
@@
Beyond the basic comparisons against DM, IPS, PI, and MIPS, we finally compare LIPS with a set of hybrid estimators including DR (Dudík et al., 2011), PI-DR (Su et al., 2020a), and OffCEM (Saito et al., 2023) (Appendix A defines these estimators in detail with math notations). Figures 8 and 8 show this comparison on Wiki10-31K and Eurlex-4K for varying slate sizes $(L\in\{4,6,\ldots,12\})$. Figures 8 and 8 compare the methods on the datasets for varying data sizes ($n\in\{1000,2000,\ldots,16000\}$). The results demonstrate that LIPS outperforms these hybrid estimators for a range of experiment values ($L$ and $n$) and non-linear reward functions. This is because DR has the variance issue due to its slate-wise importance weighting while PI-DR suffers from bias due to its linearity assumption and variance due to its slot-wise importance weighting. OffCEM is an extension of MIPS and uses a sufficient abstraction, but it is not optimized towards the MSE, and thus produces much larger variance than LIPS. These empirical observations suggest that, for OPE of slate bandits, tuning the definition of importance weights by optimizing slate abstractions with our method is more crucial and effective than introducing a reward estimator ($\hat{q}$) as in the hybrid estimators.